# FiniteTimeBibliography.bib

### v1

@@ D.5 Energy shell thermalization from quantum dynamical echoes
$\operatorname{Tr}\left[\hat{\gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]\geq\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d},\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\widetilde{w}_{+}(E_{n}-E_{m})\widetilde{v}_{+}\left(\frac{E_{n}+E_{m}}{2}-E_{c}\right)\left\lvert\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert^{2}.$ (238)
@@
$\frac{1}{d}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d},\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\left\lvert\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert^{2}<\frac{\operatorname{Tr}[\hat{\Pi}_{A}]}{WVd}\operatorname{Tr}\left[\hat{\gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right],$ (239)
@@
$\operatorname{Tr}\left[\hat{\gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]=\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{n,m}\widetilde{w}(E_{n}-E_{m})\widetilde{v}\left(\frac{E_{n}+E_{m}}{2}-E_{c}\right)\left\lvert\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert^{2}$ (240)
@@
$(\text{Terms with }\lvert E_{n}-E_{m}\rvert<\Delta E\text{ and }E_{n},E_{m}\in\Sigma_{d})$ (241)
$+$ (Terms with |$(\text{Terms with }\lvert E_{n}-E_{m}\rvert<\Delta E\text{ and }[(E_{n}\notin\Sigma_{d})\text{ or }(E_{m}\notin\Sigma_{d})])$ (242)
$+$ (Terms with |$(\text{Terms with }\lvert E_{n}-E_{m}\rvert\geq\Delta E\text{ and}(E_{n}+E_{m})/2\notin\mathcal{E}_{\Delta E})$ (243)
$+$ (Terms with |$(\text{Terms with }\lvert E_{n}-E_{m}\rvert\geq\Delta E\text{ and}(E_{n}+E_{m})/2\in\mathcal{E}_{\Delta E})$ (244)
For the first kind of terms in Line (241), using the triangle inequality and $\lvert\widetilde{v}(E)\rvert\leq 1$ and $\lvert\widetilde{w}(\delta E)\rvert\leq 1$ (from positivity, but not necessarily complete positivity), we have:
@@
For the second kind of terms in Line (242), we use the triangle inequality with $\lvert\widetilde{v}(E)\rvert<v_{0}$ and $\lvert\widetilde{w}(\delta E)\rvert\leq 1$; the former condition on $v$ follows by assumption, due to the fact that if either of $E_{n},E_{m}$ are not in $\Sigma_{d}$ but are subject to $\lvert E_{n}-E_{m}\rvert<\Delta E$, then the interval $[(E_{n}+E_{m})/2-\Delta E,(E_{n}+E_{m})/2+\Delta E]$ is not contained in $\mathcal{E}$, and therefore $(E_{n}+E_{m})/2\notin\mathcal{E}_{\Delta E}$. In addition, we can add nonnegative terms with $v_{0}/\operatorname{Tr}[\hat{\Pi}_{A}]$ multiplying the squared magnitude of the matrix elements of $(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}})$ within the energy shell to obtain the trace of its square (similar to “completing the square”), and the result must be greater than or equal to the original expression due to adding nonnegative terms. This gives the inequality:
@@
$\operatorname{Tr}\left[\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)^{2}\right]\leq\operatorname{Tr}[\hat{\Pi}_{A}^{2}]=\operatorname{Tr}[\hat{\Pi}_{A}].$ (247)
Applying a similar strategy to Line (243), where in addition to $\lvert\widetilde{v}(E)\rvert<v_{0}$ as $(E_{n}+E_{m})/2\notin\mathcal{E}_{\Delta E}$, we use $\lvert\widetilde{w}(\delta E)\rvert\leq w_{0}$ as $\lvert E_{n}-E_{m}\rvert<\Delta E$ and add the remaining absolute squared matrix elements of $(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}})$ multiplied by $w_{0}v_{0}/\operatorname{Tr}[\hat{\Pi}_{A}]$ to get:
@@
Finally, for Line (244), we use $\lvert\widetilde{w}(\delta E)\rvert\leq w_{0}$ and $\lvert\widetilde{v}(E)\rvert\leq 1$ and apply a similar strategy as above to obtain the trace of the squared operator to get
@@
$\operatorname{Tr}\left[\hat{\gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]<\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\operatorname{Tr}\left\{\left[\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right]^{2}_{\Sigma_{d},\Delta E}\right\}+v_{0}+v_{0}w_{0}+w_{0},$ (250)

### v2

@@ D.5 Energy shell thermalization from quantum dynamical echoes
$\operatorname{Tr}\left[\hat{\gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]\geq\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{\begin{subarray}{c}n,m:\\ \lvert E_{n}\rangle,\lvert E_{m}\rangle\in\Sigma_{d},\\ \lvert E_{n}-E_{m}\rvert<\Delta E\end{subarray}}\widetilde{w}_{+}(E_{n}-E_{m})\widetilde{v}_{+}\left(\frac{E_{n}+E_{m}}{2}-E_{c}\right)\left\lvert\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert^{2}.$ (D.43)
@@
[…]
@@
$\operatorname{Tr}\left[\hat{\gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]=\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\sum_{n,m}\widetilde{w}(E_{n}-E_{m})\widetilde{v}\left(\frac{E_{n}+E_{m}}{2}-E_{c}\right)\left\lvert\langle E_{m}\rvert\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\lvert E_{n}\rangle\right\rvert^{2}$ (D.45)
@@
$(\text{Full expression})=$
$(\text{Terms with }\lvert E_{n}-E_{m}\rvert<\Delta E\text{ and }E_{n},E_{m}\in\Sigma_{d})$ (D.46)
$+$ $(\text{Terms with }\lvert E_{n}-E_{m}\rvert<\Delta E\text{ and }[(E_{n}\notin\Sigma_{d})\text{ or }(E_{m}\notin\Sigma_{d})])$ (D.47)
$+$ $(\text{Terms with }\lvert E_{n}-E_{m}\rvert\geq\Delta E\text{ and }(E_{n}+E_{m})/2\notin\mathcal{E}_{\Delta E})$ (D.48)
$+$ $(\text{Terms with }\lvert E_{n}-E_{m}\rvert\geq\Delta E\text{ and }(E_{n}+E_{m})/2\in\mathcal{E}_{\Delta E})$ (D.49)
For the first kind of terms in Line (D.46), using the triangle inequality and $\lvert\widetilde{v}(E)\rvert\leq 1$ and $\lvert\widetilde{w}(\delta E)\rvert\leq 1$ (from positivity, but not necessarily complete positivity), we have:
@@
$\operatorname{Tr}\left[\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)^{2}\right]\leq\operatorname{Tr}[\hat{\Pi}_{A}^{2}]=\operatorname{Tr}[\hat{\Pi}_{A}].$ (D.52)
@@
[…]
Finally, for Line (D.49), we use $\lvert\widetilde{w}(\delta E)\rvert\leq w_{0}$ and $\lvert\widetilde{v}(E)\rvert\leq 1$ and apply a similar strategy as above to obtain the trace of the squared operator to get
@@
$\operatorname{Tr}\left[\hat{\gamma}_{A}^{\text{shell}}\left(\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right)\right]<\frac{1}{\operatorname{Tr}[\hat{\Pi}_{A}]}\operatorname{Tr}\left\{\left[\hat{\Pi}_{A}-\langle\hat{\Pi}_{A}\rangle_{\Sigma_{d}}\hat{\mathds{1}}\right]^{2}_{\Sigma_{d},\Delta E}\right\}+v_{0}+v_{0}w_{0}+w_{0},$ (D.55)